# COLT: Towards Completeness-Oriented Tool Retrieval for Large Language Models

### v1

@@ Abstract.
Recently, the integration of external tools with Large Language Models (LLMs) has emerged as a promising approach to overcome the inherent constraints of their pre-training data. However, real-world applications often involve a diverse range of tools, making it infeasible to incorporate all tools directly into LLMs due to constraints on input length and response time. Therefore, to fully exploit the potential of tool-augmented LLMs, it is crucial to develop an effective tool retrieval system. Existing tool retrieval methods techniques mainly rely on semantic matching between user queries and tool descriptions, which often results in the selection of redundant tools. As a result, these methods fail to provide a complete set of diverse tools necessary for addressing the multifaceted problems encountered by LLMs. In this paper, we propose a novel model-agnostic COllaborative Learning-based Tool Retrieval approach, COLT, which captures not only the semantic similarities between user queries and tool descriptions but also takes into account the collaborative information of tools. Specifically, we first fine-tune the PLM-based retrieval models to capture the semantic relationships between queries and tools in the semantic learning stage. Subsequently, we construct three bipartite graphs among queries, scenes, and tools and introduce a dual-view graph collaborative learning framework to capture the intricate collaborative relationships among tools during the collaborative learning stage. Extensive experiments on both the open benchmark and the newly introduced ToolLens dataset show that COLT achieves superior performance. Notably, the performance of BERT-mini (11M) with our proposed model framework outperforms BERT-large (340M), which has 30 times more parameters. Additionally, we plan to publicly release the ToolLens dataset to support further research in tool retrieval.
@@ 1. Introduction
Recently, large language models (LLMs) have demonstrated remarkable progress across various natural language processing tasks (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2023; Touvron et al., 2023; Dai et al., 2023b). However, they often struggle with solving highly complex problems and providing up-to-date knowledge due to the constraints of their pre-training data (Mallen et al., 2022; Vu et al., 2023). A promising approach to overcome these limitations is tool learning (Schick et al., 2023; Parisi et al., 2022; Li et al., 2023; Ye et al., 2024), which enables LLMs to dynamically interact with external tools, significantly facilitating access to real-time data and the execution of complex computations. By integrating tool learning, LLMs transcend the confines of their outdated or limited pre-trained knowledge (Brown et al., 2020), offering responses to user queries with significantly improved accuracy and relevance (Huang et al., 2023; Qin et al., 2023b). However, real-world systems typically involve a large number of tools, making it impractical to take the descriptions of all tools as input for LLMs due to length limitations and latency constraints. Thus, as illustrated in Figure 1(a), developing an effective tool retrieval system becomes essential to fully exploit the potential of tool-augmented LLMs (Gao et al., 2024).
Typically, existing tool retrieval approaches directly employ dense retrieval techniques (Qin et al., 2023b; Gao et al., 2024; Yuan et al., 2024), solely focusing on matching semantic similarities between queries and tool descriptions. However, these approaches may fall short when addressing multifaceted queries that require a collaborative effort from multiple tools to formulate a comprehensive response. For instance, in Figure 1(b), consider a user’s request to calculate the value of 5 ounces of gold plus 1 million AMZN stocks in CNY. Such a query requires the simultaneous use of tools for gold prices, stock values, and currency exchange rates. The absence of any of these tools yields an incomplete answer 111The complete responses are provided in Appendix C.. In this example, dense retrieval methods that rely solely on semantic matching may retrieve multiple tools related to stock prices while neglecting others. This highlights a significant limitation of dense retrieval methods that overlook the necessity for tools to interact collaboratively. Thus, ensuring the completeness of retrieved tools is an essential aspect of a tool retrieval system, which is often neglected by traditional retrieval approaches.
@@ 3.3. Semantic Learning
where and denote the mean pooling vectors from the final layer of the selected PLM, and represents the cosine similarity function.
@@ 3.4. Collaborative Learning
In summary, our dual-view graph collaborative learning framework yields two sets of embeddings: and from the scene-centric view, and and from the tool-centric view, for queries and scenes, respectively. Then, the final matching score of each given query-tool pair is implemented according to the following formula:
@@
As shown in Figure 2 (d), to effectively capture high-order collaborative relationships between tools and align the cooperative interactions across two views, we utilize a cross-view contrastive loss. Specifically, the representations of queries and scenes can be learned by optimizing the cross-view InfoNCE (Gutmann and Hyvärinen, 2010) loss:
@@
Therefore, the list-wise multi-label loss function is formulated to minimize the discrepancy between these two probability distributions:
@@ 4. Datasets
ToolBench. ToolBench (Qin et al., 2023b) is a benchmark commonly used to evaluate the capability of LLMs in tool usage. In our experiments, we notice that its three subsets exhibit distinct characteristics. The first subset (I1) focuses on single-tool scenarios, which diverges from our emphasis on multi-tool tasks. In contrast, the second subset (I2) creates queries by selecting tools within the same category, and the third subset (I3) forms queries by selecting tools from a wider range. These characteristics of I2 and I3 align with our focus on multi-tool tasks, leading us to select them as the primary datasets for our experiments.
@@
As shown in Figure 3, the creation of ToolLens involves a novel five-step methodology 222The prompts designed for GPT-4 to construct proposed ToolLens are provided in Appendix A: 1) Tool Selection: To create a high-quality tool dataset, we initially conduct rigorous filtering the original tool collection from ToolBench, focusing on tools that are both available and directly callable. To better align our dataset with the practical query needs of real-world users, we exclude tools that are not relevant to everyday user queries, such as those for authentication or testing, retaining 464 high-quality, callable tools; 2) Scene Mining: We design an instruction to prompt GPT-4 to generate potential scenes scenes that are relevant to the detailed descriptions of the selected tools, and ensure their validity through human verification. 3) Query Generation: We then design another instruction to employ GPT-4 to generate queries based on the provided scene and the parameters required for tool calling. Notably, we avoid providing the complete tool description to GPT-4 to avoid the generated query being closely aligned with the tool. 4) Tool Aggregation: The queries generated in the aforementioned way are only relevant to a single tool. To enhance the relevance of queries across multiple tools, we reprocess them through GPT-4 to identify categories of tools that could be relevant, which are then aligned with our tool set through dense retrieval and manual verification. 5) Query Rewriting: Finally, we utilize GPT-4 to revise queries to incorporate all necessary parameters by providing it with both the initial query and a list of required parameters, thereby yielding concise yet intentionally multifaceted queries that better mimic real-world user interactions. It is worth noting that we incorporate a human verification process at each step to ensure the quality of the generated data.
This comprehensive construction pipeline ensures that ToolLens accurately simulates the dynamics of real-world tool retrieval scenarios. We develop the ToolLens dataset using this method, which includes 18,770 queries and 464 tools, with each query being associated with verified tools.
@@
Discussion and Quality Verification. Different from prior datasets for tool learning, ToolLens uniquely focuses on creating queries that are natural, concise, and intentionally multifaceted to more accurately reflect the complex demands in real-world scenarios. To assess the quality of ToolLens, following previous works (Gao et al., 2024; Liu et al., 2023a; Sottana et al., 2023), we employ GPT-4 as an evaluator and human evaluation where three well-educated doctor students are invited to evaluate 50 randomly sampled cases from ToolLens, ToolBench and TOOLE in the following two aspects:(1) Natural-query: whether the query is natural. (2) Multifaceted intentions: whether the user intent is multifaceted. The results are illustrated in Table 2. In most cases, ToolLens outperforms ToolBench and TOOLE 333Some cases of comparison between ToolLens and ToolBench are provided in Appendix B.. Furthermore, using GPT-4 as the evaluator shows a high degree of consistency with human evaluation trends, which underscores the validity of employing GPT-4 as an evaluator.
@@ 5. Experiments
In this section, we first describe the experimental setups and then conduct an extensive evaluation and analysis of the proposed COLT.
@@ 5.1. Experimental Setups
where denotes the set of ground-truth tools for query , represents the top- tools retrieved for query , and is an indicator function that returns 1 if the retrieval results include all ground-truth tools within the top- results for query , and 0 otherwise.
@@
We utilize the BEIR (Thakur et al., 2021) framework to implement the dense retrieval baselines, set the training epochs to with the learning rate of , weight decay of , and using the AdamW optimizer. As our approach is model-agnostic, we directly apply dense retrieval for the semantic learning stage, and in the collaborative learning stage, we set the batch size as 2048 and carefully tune the learning rate among , the weight decay among , as well as the layer number among . All the experiments are conducted on NVIDIA RTX A6000 48G GPUs.
@@ 5.2. Experimental Results
Table 3 presents the overall results of different tool retrieval methods on ToolLens, ToolBench (I2), and ToolBench (I3). From the results, we have the following observations and conclusions:
We can observe that the performance of traditional dense retrieval models in zero-shot scenarios is quite poor, even inferior to that of BM25. This indicates that these models may not be well-suited for tool retrieval tasks. Conversely, the BM25 model significantly lags behind fine-tuned PLM-based dense retrieval methods, underscoring the superior capability of the latter in leveraging contextual information for more effective tool retrieval. Despite this advantage, PLM-based methods fall short in the COMP metric, which is specifically designed for evaluating completeness in tool retrieval scenarios. This suggests that while effective for general retrieval tasks, PLM-based methods may not fully meet the unique demands of tool retrieval.
@@
To verify that improvements of COLT in tool retrieval truly enhance downstream real-world tool learning applications, we further employ the pairwise comparison method to conduct a validation study (Dai et al., 2023a; Sun et al., 2023; Liang et al., 2023). We randomly select 100 queries from the test set of ToolLens and use various retrieval models to retrieve the top-3 tools for each query. Then we utilize GPT-4 as an evaluator, examining the responses generated with different retrieved tools across four dimensions: Coherence, Relevance, Comprehensiveness, and Overall. Specifically, the user query and a pair of responses are utilized as prompts to guide GPT-4 in determining the superior response. Additionally, we also consider that LLMs may respond differently to the order in which text is presented in the prompt (Lu et al., 2022; Tang et al., 2023b; Hou et al., 2024; Liu et al., 2023b). To mitigate potential biases associated with this, we conduct each comparison twice, reversing the response order for the second evaluation, which ensures a more reliable assessment.
We establish a tournament-style competition where models compete against each other using the Elo ratings system, which is widely employed in chess and other two-player games to measure the relative skill levels of the players (Dettmers et al., 2023; Wu and Aji, 2023). Each player is assigned an Elo score. Given two players and with their Elo ratings and , the expected scores and for these two players are:
The outcome of matches between players leads to adjustments in their Elo scores, and the updated Elo ratings and are:
where and are the actual scores of players and , is the -factor that determines the maximum amount of points a player’s rating can change from a single game outcome.
@@
Following previous works (Chiang et al., 2023), we start with a score of and set . Additionally, to minimize the impact of match sequences on Elo score calculations, we conduct these computations times using various random seeds to control for ordering effects.
The results in Table 4 show that superior tool retrieval models can significantly improve downstream tool learning performance. Moreover, responses generated with the tools retrieved from COLT notably outperform those from other methods, achieving the highest Elo ratings in all four assessed dimensions. These results highlight the pivotal role of effective tool retrieval in enhancing the performance of downstream applications and further confirm the superiority of COLT.
@@ 5.3. Further Analysis
To verify the adaptability and effectiveness of COLT across varying sizes of PLMs, we explore its integration with a range of BERT models, from BERT-mini to BERT-large. This analysis aims to determine whether COLT could generally enhance tool retrieval performance across different model sizes. Figure 4 presents the results, illustrating a clear trend: while the performance of the base model naturally improves with larger PLM sizes, the integration of COLT consistently boosts performance across all sizes. Remarkably, even BERT-mini equipped with COLT, significantly outperforms a much larger BERT-large model (30x larger) operating without our COLT. These results underscore the generalization and robustness of COLT, demonstrating its potential to significantly improve tool retrieval performance for PLMs of any scale.

### v2

# Towards Completeness-Oriented Tool Retrieval for Large Language Models

@@ Abstract.
Recently, integrating external tools with Large Language Models (LLMs) has gained significant attention as an effective strategy to mitigate the limitations inherent in their pre-training data. However, real-world systems often incorporate a wide array of tools, making it impractical to input all tools into LLMs due to length limitations and latency constraints. Therefore, to fully exploit the potential of tool-augmented LLMs, it is crucial to develop an effective tool retrieval system. Existing tool retrieval methods primarily focus on semantic matching between user queries and tool descriptions, frequently leading to the retrieval of redundant, similar tools. Consequently, these methods fail to provide a complete set of diverse tools necessary for addressing the multifaceted problems encountered by LLMs. In this paper, we propose a novel model-agnostic COllaborative Learning-based Tool Retrieval approach, COLT, which captures not only the semantic similarities between user queries and tool descriptions but also takes into account the collaborative information of tools. Specifically, we first fine-tune the PLM-based retrieval models to capture the semantic relationships between queries and tools in the semantic learning stage. Subsequently, we construct three bipartite graphs among queries, scenes, and tools and introduce a dual-view graph collaborative learning framework to capture the intricate collaborative relationships among tools during the collaborative learning stage. Extensive experiments on both the open benchmark and the newly introduced ToolLens dataset show that COLT achieves superior performance. Notably, the performance of BERT-mini (11M) with our proposed model framework outperforms BERT-large (340M), which has 30 times more parameters. Furthermore, we will release ToolLens publicly to facilitate future research on tool retrieval.
@@ 1. Introduction
Recently, large language models (LLMs) have demonstrated remarkable progress across various natural language processing tasks (Brown et al., 2020; Chowdhery et al., 2023; Achiam et al., 2023; Touvron et al., 2023). However, they often struggle with solving highly complex problems and providing up-to-date knowledge due to the constraints of their pre-training data (Mallen et al., 2022; Vu et al., 2023). A promising approach to overcome these limitations is tool learning (Schick et al., 2023; Parisi et al., 2022; Li et al., 2023; Ye et al., 2024; Qu et al., 2024), which enables LLMs to dynamically interact with external tools, significantly facilitating access to real-time data and the execution of complex computations. By integrating tool learning, LLMs transcend the confines of their outdated or limited pre-trained knowledge (Brown et al., 2020), offering responses to user queries with significantly improved accuracy and relevance (Huang et al., 2023; Qin et al., 2023b). However, real-world systems typically involve a large number of tools, making it impractical to take the descriptions of all tools as input for LLMs due to length limitations and latency constraints. Thus, as illustrated in Figure 1(a), developing an effective tool retrieval system becomes essential to fully exploit the potential of tool-augmented LLMs (Gao et al., 2024).
Typically, existing tool retrieval approaches directly employ dense retrieval techniques (Qin et al., 2023b; Gao et al., 2024; Yuan et al., 2024), solely focusing on matching semantic similarities between queries and tool descriptions. However, these approaches may fall short when addressing multifaceted queries that require a collaborative effort from multiple tools to formulate a comprehensive response. For instance, in Figure 1(b), consider a user’s request to calculate the value of 5 ounces of gold plus 1 million AMZN stocks in CNY. Such a query requires the simultaneous use of tools for gold prices, stock values, and currency exchange rates. The absence of any of these tools yields an incomplete answer. In this example, dense retrieval methods that rely solely on semantic matching may retrieve multiple tools related to stock prices while neglecting others. This highlights a significant limitation of dense retrieval methods that overlook the necessity for tools to interact collaboratively. Thus, ensuring the completeness of retrieved tools is an essential aspect of a tool retrieval system, which is often neglected by traditional retrieval approaches.
@@ 3.3. Semantic Learning
where and are the mean pooling vectors from the final layer of the PLM, and represents the cosine similarity function.
@@ 3.4. Collaborative Learning
In summary, our dual-view graph collaborative learning framework yields two sets of embeddings: and from the scene-centric view, and and from the tool-centric view for queries and scenes, respectively. Then, the final matching score of each query-tool pair is calculated using the following formula:
@@
As shown in Figure 2 (d), we capture high-order collaborative relationships between tools and align the cooperative interactions across two views using a cross-view contrastive loss. Specifically, the representations of queries and scenes can be learned by optimizing the cross-view InfoNCE (Gutmann and Hyvärinen, 2010; Tang et al., 2024) loss:
@@
Therefore, the list-wise multi-label loss function minimizes the discrepancy between these two probability distributions:
@@ 4. Datasets
ToolBench. ToolBench (Qin et al., 2023b) is a benchmark commonly used to evaluate the capability of LLMs in tool usage. In our experiments, we notice that its three subsets exhibit distinct characteristics. The first subset (I1) focuses on single-tool scenarios, which diverges from our emphasis on multi-tool tasks. However, both the second subset (I2) and the third subset (I3) align with our focus on multi-tool tasks. Therefore, we chose I2 and I3 as the primary datasets for our experiments.
@@
As shown in Figure 3, the creation of ToolLens involves a novel five-step methodology: 1) Tool Selection: To create a high-quality tool dataset, we rigorously filter ToolBench, focusing on 464 available and directly callable tools relevant to everyday user queries, excluding those for authentication or testing. 2) Scene Mining: We prompt GPT-4 to generate potential scenes that are relevant to the detailed descriptions of the selected tools, and ensure their validity through human verification. 3) Query Generation: We then employ GPT-4 to generate queries based on the provided scene and the parameters required for tool calling. Notably, we avoid providing the complete tool description to GPT-4 to avoid the generated query being closely aligned with the tool. 4) Tool Aggregation: The queries generated in the aforementioned way are only relevant to a single tool. To enhance the relevance of queries across multiple tools, we reprocess them through GPT-4 to identify categories of tools that could be relevant, which are then aligned with our tool set through dense retrieval and manual verification. 5) Query Rewriting: Finally, we utilize GPT-4 to revise queries to incorporate all necessary parameters by providing it with both the initial query and a list of required parameters, thereby yielding concise yet intentionally multifaceted queries that better mimic real-world user interactions. It is worth noting that we incorporate a human verification process at each step to ensure data quality.
This comprehensive construction pipeline ensures ToolLens accurately simulates real-world tool retrieval dynamics. The resulting ToolLens dataset includes 18,770 queries and 464 tools, with each query being associated with verified tools.
@@
Discussion and Quality Verification. Unlike prior datasets, ToolLens uniquely focuses on creating natural, concise, and multifaceted queries to reflect real-world demands. To assess the quality of ToolLens, following previous works (Gao et al., 2024; Liu et al., 2023a; Sottana et al., 2023), we employ GPT-4 as an evaluator and human evaluation where three well-educated doctor students are invited to evaluate 50 randomly sampled cases from ToolLens, ToolBench and TOOLE in the following two aspects:(1) Natural-query: whether the query is natural. (2) Multifaceted intentions: whether the user intent is multifaceted. The results are illustrated in Table 2. In most cases, ToolLens outperforms ToolBench and TOOLE. Furthermore, using GPT-4 as the evaluator shows a high degree of consistency with human evaluation trends, which underscores the validity of employing GPT-4 as an evaluator.
@@ 5. Experiments
In this section, we first describe the experimental setups and then conduct an extensive evaluation and analysis of the proposed COLT. The source code and the proposed ToolLens dataset are publicly available at https://github.com/quchangle1/COLT.
@@ 5.1. Experimental Setups
where is the set of ground-truth tools for query , represents the top- tools retrieved for query , and is an indicator function that returns 1 if the retrieval results include all ground-truth tools within the top- results for query , and 0 otherwise.
@@
We utilize the BEIR (Thakur et al., 2021) framework for dense retrieval baselines, setting the training epochs to with the learning rate of , weight decay of , and using the AdamW optimizer. Our model-agnostic approach directly applies dense retrieval for semantic learning. During collaborative learning, we set the batch size as 2048 and carefully tune the learning rate among , the weight decay among , as well as the layer number among .
@@ 5.2. Experimental Results
Table 3 presents the results of different tool retrieval methods on ToolLens, ToolBench (I2 and I3). From the results, we have the following observations and conclusions:
We can observe that traditional dense retrieval models perform poorly in zero-shot scenarios, even inferior to that of BM25. This indicates that these models may not be well-suited for tool retrieval tasks. Conversely, the BM25 model significantly lags behind fine-tuned PLM-based dense retrieval methods, underscoring the superior capability of the latter in leveraging contextual information for more effective tool retrieval. Despite this advantage, PLM-based methods fall short in the COMP metric, which is specifically designed for evaluating completeness in tool retrieval scenarios. This suggests that while effective for general retrieval tasks, PLM-based methods may not fully meet the unique demands of tool retrieval.
@@
To verify that improvements of COLT in tool retrieval truly enhance downstream tool learning applications, we conduct a validation study using the pairwise comparison method (Dai et al., 2023; Sun et al., 2023; Liang et al., 2023). We randomly select 100 queries from the test set of ToolLens and use various retrieval models to retrieve the top-3 tools for each query. Then we utilize GPT-4 as an evaluator, examining the responses generated with different retrieved tools across four dimensions: Coherence, Relevance, Comprehensiveness, and Overall. Specifically, the user query and a pair of responses are utilized as prompts to guide GPT-4 in determining the superior response. Additionally, we also consider that LLMs may respond differently to the order in which text is presented in the prompt (Lu et al., 2022; Tang et al., 2023b; Hou et al., 2024; Liu et al., 2023b). So each comparison is conducted twice with reversed response order to mitigate potential biases from text order, ensuring a more reliable assessment.
@@
We establish a tournament-style competition using the Elo ratings system, which is widely employed in chess and other two-player games to measure the relative skill levels of the players (Dettmers et al., 2023; Wu and Aji, 2023). Following previous works (Chiang et al., 2023), we start with a score of and set -factor to . Additionally, to minimize the impact of match sequences on Elo scores, we conduct these computations times using various random seeds to control for ordering effects.
The results in Table 4 show that superior tool retrieval models can significantly improve downstream tool learning performance. Moreover, responses generated with the tools retrieved from COLT notably outperform those from other methods, achieving the highest Elo ratings in all four assessed dimensions. These results highlight the pivotal role of effective tool retrieval in tool learning applications and further confirm the superiority of COLT.
@@ 5.3. Further Analysis
To verify the adaptability and effectiveness of COLT across varying sizes of PLMs, we explore its integration with a range of BERT models, from BERT-mini to BERT-large. This analysis aims to determine whether COLT could generally enhance tool retrieval performance across different model sizes. Figure 4 shows that while the performance of the base model naturally improves with larger PLM sizes, the integration of COLT consistently boosts performance across all sizes. Remarkably, even BERT-mini equipped with COLT, significantly outperforms a much larger BERT-large model (30x larger) operating without our COLT. These results underscore the generalization and robustness of COLT, demonstrating its potential to significantly improve tool retrieval performance for PLMs of any scale.